Suppose that a random variable has the normal distribution N(6, 4). Find the
expectation of X^2, E(X^2).
1
Expert's answer
2019-11-26T09:12:31-0500

We have "\mu=6, \sigma^2=4"

We know that "Var(X)=E(X^2)-(E(X))^2"

From that , "E(X^2)=Var(X)+(E(X))^2"

"E(X^2)=\sigma^2+\mu^2"

"=4+6^2"

"=4+36=40"